OpenStudy (anonymous):

A spherical balloon is being inflated at the rate of 4 cubic ft/min. The rate, in square feet/min, at which the surface area is increasing when the volume is 32pi/3 ft^3 is...?

OpenStudy (anonymous):

When we are inflating a balloon, we are just pumping volume of air into it. .'.dV/dt = 4cu ft/min V = (4/3)πr^3 .'.dV/dt = (4/3)π[3r^2(dr/dt)] or, 4 = (4/3)π[3r^2(dr/dt)] or, (1/π) = r^2(dr/dt).............(1) When, volume is 32π/3 cu ft (4/3)πr^3 = 32π/3 r^3 = 8 or, r=2. .'.dr/dt = 1/(4π) Now, surface area = S = 4πr^2 .'.dS/dt(at r=2) = 8πr(dr/dt) = 8π*2*[1/(4π)] = 16/4 = 4 ft^2/min .'.rate of increase of surface area of a sphere = 4ft^2 / min

OpenStudy (anonymous):

answer choices: a. 4pi b. 2 c. 4 d. 1 e. 2pi I know this is a related rates problem and I just cant find a good example in my book or notes! I would appreciate any help!

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Oh wow! Thanks for such a concise and quick reply! I looked through your work and it looks great! I understand this now, and I'll work through it one more time just to solidify it for myself. Again thank you, and I'll award you Best Response!

OpenStudy (anonymous):

you welcome I just like helping people lol

OpenStudy (anonymous):

Well thanks much then! I'll close this question.